# EFRame: Deeper Reasoning via Exploration-Filter-Replay Reinforcement Learning Framework

### v3

@@ 5.1 Experimental setup
Training Datasets. Our experiments are conducted on two datasets to evaluate the effectiveness of the algorithm in both in-domain and out-of-domain settings: (1) Geometry3K focused on geometric problems Lu et al. (2021a), (2) a 6k out-of-domain dataset randomly sampled from the dataset proposed in  Wei et al. (2025b; a), which randomly sampled from established open-source resources, including Geometry3K Lu et al. (2021a), GeoQA Chen et al. (2021), GeoQA-Plus Cao & Xiao (2022), Geos Seo et al. (2015), AI2D Kembhavi et al. (2016), TQA Kim et al. (2018), FigureQA Kahou et al. (2017), TabMWP Lu et al. (2022b), ChartQA Masry et al. (2022), IconQA Lu et al. (2021b), Clevr-Math Lindström & Abraham (2022), M3CoT Chen et al. (2024), and ScienceQA Lu et al. (2022a).
Evaluation Benchmarks. Our experiments consist of two parts. First, we evaluate the overall performance of the algorithms on four open multimodal math benchmarks: MathVision Wang et al. (2024a), MathVista Lu et al. (2023), MathVerse Zhang et al. (2024), and We-Math Qiao et al. (2024). These benchmarks cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty. Second, we conduct a detailed analysis of algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
Baselines. We compare our models with a range of advanced multimodal large language models. These include GPT-4o Hurst et al. (2024) and Claude-3.5-Sonnet Anthropic (2024), which are state-of-the-art closed-source MLLMs. QvQ-Preview-72B[2] and Skywork R1V-32B  Peng et al. (2025) are multimodal large reasoning models specifically designed for R1-like chain-of-thought capabilities. At the 7B scale, we compare with MM-Eureka-7B Meng et al. (2025), OpenVLThinker-7B Deng et al. (2025), ThinkLite-VL-7B Wang et al. (2025b), and VLAA-Thinker-Qwen2.5VL-7B Chen et al. (2025a). All of these 7B models are based on Qwen2.5-VL and are enhanced for mathematical reasoning via fine-tuning, reinforcement learning, or iterative self-improvement techniques. In addition, we compare the performance of GRPO and DAPO on the Qwen2.5-VL-7B-Instruct model to assess the impact of different strategies under consistent model and training settings.
Implementation details. Our work is based on EasyR1 and VeRL frameworks Yaowei et al. (2025); Sheng et al. (2025). We initialize all policy models with Qwen2.5-VL-7B-Instruct and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$. GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward to avoid relative advantage bias (see detail in Appendix A), while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens. For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
Table 2: The performance comparison on the Geometry3K dataset shows that our method achieves a significant improvement over existing approaches.
Models Qwen2.5-VL-Instruct Ours GRPO DAPO
Accuracy 38.44 55.24(+16.80) 50.75 (+12.31) 49.09 (+10.65)
Time/step(s) - 442($\times$1.35) 327($\times$1.00) 182($\times$0.57)
@@ 5.2 Main results
Performance. In our experiments, we employ the Exploration-Filter-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets. As shown in Table 1, the proposed method delivers consistently strong results across four multimodal reasoning benchmarks, with an average gain of +5.66 over the base model (Qwen2.5-VL-7B-Instruct), outperforming GRPO by 19.9%. On the in-domain Geometry3K dataset, the improvement is even more pronounced: Table 2 reports a +16.80 gain, representing a 36.5% relative advantage over GRPO.
Exploration. As shown in Fig. 4, our framework maintains the policy entropy around 1, striking a balance between sufficient exploration and exploitation. In the later training stages, when the vanilla GRPO tends to suffer from vanishing advantages, our method further increases entropy to encourage continued exploration of informative reward signals.
Stability. Our framework leverages additional rollouts to enrich reward signals, while an online filtering mechanism mitigates training variance by excluding low-quality samples. Extensive experiments reveal that the proposed method circumvents collapse issues—occasionally encountered in GRPO and DAPO. Furthermore, the steadily increasing response length during training reflects the model’s progressively improved reasoning ability.
5.3 Entropy Control
To achieve a well-balanced exploration process, we incorporate entropy control into EFRame by adjusting two key hyperparameters: the additional rollout temperature $t_{a}$ and the replay buffer size $R_{s}$. Specifically, increasing the temperature in the additional rollout phase encourages broader exploration, enabling the model to discover low-probability but potentially high-reward trajectories. At the same time, we increase the $R_{s}$ to promote convergence and prevent excessive entropy growth. Together, these mechanisms allow for a controlled yet sufficiently rich exploration process—ensuring that the model explores meaningfully without destabilizing training due to overly random behavior.
@@ 5.3 Entropy Control
Accuracy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
$R_{s}=5120$ 53.74 52.41 55.24
$R_{s}=2560$ 52.75 54.41 52.75
$R_{s}=0$ 52.41 51.08 50.25
Entropy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
$R_{s}=5120$ 0.089 0.305 0.726
$R_{s}=2560$ 0.091 0.312 2.366
$R_{s}=0$ 0.092 2.437 4.288
Figure 5: We report the peak accuracy within the first 200 training steps and the average entropy over the first 150 steps. Experiments on the GEO3K dataset using Qwen2.5-VL-7B-Instruct under different settings of additional rollout temperature $t_{a}$ and replay buffer size $R_{s}$ demonstrate that increasing $t_{a}$ enhances exploration, reflected in higher entropy, while enlarging $R_{s}$ promotes convergence, leading to lower entropy. Optimal performance is achieved when exploration and convergence are well-balanced.
We observe in Fig. 5 that increasing the additional rollout temperature encourages broader exploration, leading to a rise in policy entropy. In contrast, enlarging the replay buffer size, which increases the number of high-quality samples reused during training, promotes convergence and results in decreased entropy. These two hyperparameters exhibit opposing effects on entropy dynamics, providing a practical mechanism for entropy control within EFRame. We find that model performance peaks when the entropy stabilizes around a moderate level. Fig. 5 reports the accuracy under different $t_{a}$ and $R_{s}$, highlighting the impact of entropy control of EFRame on learning efficacy. Our experiments demonstrate that entropy control can be achieved with high precision by combining high-temperature positive sampling and experience replay, which are more effective than using entropy-based reward functions.
@@ 6 Ablation Studies
To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are essential for achieving the observed improvements in stability and overall performance.
@@ 6.1 Additional Rollout
Disabling the additional rollout module exposes the model to a fundamental trade-off inherent in GRPO-based training. When the number of samples $G_{1}$ is small, the reward signal is constrained by the base model’s pass@$G_{1}$, limiting the model’s ability to solve challenging problems and suppressing the emergence of “aha moments.”
@@
High-quality samples have the potential to activate deeper reasoning abilities within the model.
As shown in Fig. 6, the model’s reward on the training set plateaus at a relatively low level, restricting further gains in test accuracy. In the early stages of training, the entropy remains low, indicating limited exploration. However, once the model hits this performance bottleneck, entropy begins to rise—yet this delayed exploration yields minimal returns, suggesting that the model struggles to escape local optima. The additional rollout module is specifically designed to break this situation, enabling targeted deep exploration.
@@ 6.2 Online Filtering
Removing the online filtering module results in significant instability and inefficiency during training. Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distort the learning signal. This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplify gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses. As a result, training becomes slower, more chaotic, and more prone to divergence.
@@
Low-quality samples from challenging prompts often mislead optimization, causing the model to converge prematurely to suboptimal solutions and impeding further learning.
As shown in Fig. 6, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set but yields little improvement on the test set. This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability. Moreover, the forward and backward computation of these negative samples incurs substantial overhead, reducing the sample efficiency of the overall training process. Online filtering plays a crucial role in maintaining gradient clarity and training stability by selectively preserving informative samples and discarding neutral or low-impact ones.
6.3 Replay Buffer
When the replay buffer is disabled, we observe a drastic drop in performance, even falling below that of the vanilla GRPO baseline, as shown in Fig. 6. This counterintuitive result reveals that the abundant high-quality samples generated during the additional rollout phase can actually become detrimental without a proper replay mechanism.
@@ 6.3 Replay Buffer
While high-quality samples may induce high-entropy exploration, incorporating experience replay enables the model to retain and consolidate useful gradients, thus restoring training stability.
As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy. Due to their high advantage values and low sampling probabilities, high-quality samples are more likely to have their gradients clipped during training. In the absence of a replay buffer, such valuable samples contribute only short-lived and unstable gradients that are quickly overwritten. Instead of reinforcing beneficial behavior, this transient gradient signal increases policy uncertainty, pushing the model into a state of chaotic, high-entropy exploration, ultimately harming both learning stability and performance. To address this issue, we leverage experience replay to ensure that high-quality samples continuously provide effective gradients, enabling the model to converge along the new trajectory.
7 Conclusion and Discussion
In this work, we investigate the impact of sample quality on RL for large language models, with a particular focus on addressing challenging prompts. Our findings show that low-quality incorrect samples—those with low absolute advantage—tend to trap the model in local optima and offer little or even negative contribution to training. In contrast, high-quality samples with large advantages play a critical role in guiding the model through a meaningful transition from entropy-driven exploration to entropy-reducing convergence. This process is essential for eliciting deeper reasoning capabilities within the model. To this end, we propose an Exploration-Filter-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples. Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations. This strategy not only helps transform isolated “aha moments” into a sustained “aha state” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.
@@ 7 Conclusion and Discussion
Appendix A Relative Advantage Bias of GRPO
In GRPO, computing relative advantages within a group can introduce bias, especially when the reward becomes multi-dimensional. For example, as shown in Fig. 7, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e. when all responses are either correct or incorrect. In such cases, auxiliary rewards (e.g., length) dominate the advantage computation. This may result in incorrect responses being rewarded simply for being shorter, or correct responses being penalized for being longer. Such misleading gradients introduce irreversible noise into training and lead to reward hacking. Therefore, in the training involved in this paper, we use binary rewards whenever possible.
Figure 7: During training with GRPO, when the primary reward signal (accuracy) vanishes, other auxiliary rewards become dominant, which introduces bias.

### v5

@@ 4.1 Experimental setup
Training Datasets. Our experiments are conducted on three datasets to evaluate the effectiveness of the algorithm across different domains and tasks: (1) DAPO-17K, a large-scale mathematics out-of-domain (ood) training dataset designed to evaluate the algorithm’s performance on large language models (LLMs); (2) Multimodal ood dataset of size 6k randomly sampled from the dataset proposed in Wei et al. (2025b; a), which is collected from established open-source resources, including Geometry3K Lu et al. (2021a), GeoQA Chen et al. (2021), GeoQA-Plus Cao & Xiao (2022), Geos Seo et al. (2015), AI2D Kembhavi et al. (2016), TQA Kim et al. (2018), FigureQA Kahou et al. (2017), TabMWP Lu et al. (2022b), ChartQA Masry et al. (2022), IconQA Lu et al. (2021b), Clevr-Math Lindström & Abraham (2022), M3CoT Chen et al. (2024), and ScienceQA Lu et al. (2022a); (3) Geometry3K, focused on geometric problems Lu et al. (2021a), which is used to evaluate the algorithm’s in-domain performance and conduct ablation studies.
Evaluation Benchmarks. First, we evaluate the algorithm on four text-only reasoning benchmarks—AIME’24 HuggingFaceH4 (2025), MATH Lightman et al. (2023), GSM8K Cobbe et al. (2021), and Minerva Lewkowycz et al. (2022)—which are widely used to assess the problem-solving and reasoning capabilities of large language models (LLMs) across different levels of mathematical difficulty and domain coverage. These benchmarks span competition-level problems (AIME’24), advanced mathematics problem sets (MATH), grade-school word problems (GSM8K), and science/mathematics questions from Minerva. Second, we evaluate on four open multimodal benchmarks: MathVision Wang et al. (2024a), MathVista Lu et al. (2023), MathVerse Zhang et al. (2024), and We-Math Qiao et al. (2024), which cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty. Finally, we conduct a detailed analysis of the algorithms on the Geometry3K dataset to examine in-domain performance and perform ablation studies.
@@ 4.2 Main results
Models Qwen2.5-VL-7B-Instruct EFRame GRPO DAPO
Accuracy 38.44 55.41(+16.97) 50.75 (+12.31) 49.09 (+10.65)
@@
Performance. Our Exploration-Filter-Replay framework achieves substantial improvements in both in-domain and out-of-domain settings for text-only and multimodal reasoning tasks. As shown in Table 1, on the text benchmarks, our method attains an average score of 60.2, outperforming the RL baselines GRPO and DAPO by +5.1 and +2.7 average points, respectively. The largest absolute gain is observed on AIME’24 (+23.4), demonstrating the framework’s strength on challenging competition-level problems. On the multimodal benchmarks, our method reaches an average score of 55.13, outperforming GRPO by +0.94. On the in-domain Geometry3K dataset, the gain is even more pronounced: Fig. 3 reports a +16.97 improvement, corresponding to a 37.9% relative advantage over GRPO.
Exploration and Entropy Control. As shown in Fig.3, our framework maintains batch-level entropy around 1, striking a balance between exploration and exploitation. In the later stages of training, when GRPO and DAPO tend to suffer from vanishing advantages, our method keeps the exploration of informative reward signals. This is achieved by tuning two hyperparameters: the additional rollout temperature $t_{a}$ and the replay buffer size $R_{s}$. A higher $t_{a}$ broadens exploration by sampling low-probability yet high-reward trajectories, while a larger $R_{s}$ promotes convergence and suppresses excessive entropy growth. As shown in Fig.4, the two factors exert opposite influences on entropy dynamics, with optimal performance observed when entropy stabilizes at a moderate level.
Claim 1.
The combination of high-quality sampling and experience replay enables flexible and efficient entropy control.
To further improve stability, the framework leverages additional rollouts to enrich reward signals, while an online filtering mechanism reduces training variance by removing low-quality samples, resulting in more robust gradients. Extensive experiments confirm that this design effectively avoids the collapse issues observed in GRPO and DAPO.
@@
Accuracy (Entropy) $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
$R_{s}=5120$ 53.74 (0.09) 52.41 (0.31) 55.41 (0.73)
$R_{s}=2560$ 52.75 (0.09) 54.41 (0.31) 52.75 (2.37)
$R_{s}=0$ 52.41 (0.09) 51.08 (2.44) 50.25 (4.29)
Figure 4: Peak accuracy (first 200 steps) and average entropy (first 150 steps) on Geometry3K with Qwen2.5-VL-7B-Instruct under varying $t_{a}$ and $R_{s}$. Higher $t_{a}$ boosts exploration (entropy↑), larger $R_{s}$ aids convergence (entropy↓), and best performance arises from the balance state of entropy.
@@ 5 Ablation Studies
To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and experience replay—are essential for achieving the observed improvements in stability and overall performance.
@@ 5.1 Additional Rollout
[…]
@@
High-quality samples can unlock the ability of deeper reasoning in RL.
As shown in Fig. 5, training rewards plateau at low levels, constraining test accuracy. Early entropy remains low, reflecting limited exploration; once performance stalls, entropy rises but yields little benefit, indicating difficulty escaping local optima. The additional rollout module addresses this by driving targeted deep exploration.
@@ 5.2 Online Filtering
Removing the online filtering module leads to unstable and inefficient training. Without filtering, all samples are treated equally: zero-advantage ones inflate gradient variance, and low-quality negatives distort the learning signal. This effect worsens in additional rollouts, where numerous incorrect samples overwhelm rare high-quality responses, making training slower, noisier, and prone to divergence.
@@
Low-quality samples from challenging prompts mislead optimization, driving premature convergence to suboptimal solutions and preventing further learning.
As shown in Fig. 5, many negative samples initially boost training performance but provide little generalization gain, indicating faster convergence to local optima. Their forward and backward computation also adds overhead, lowering sample efficiency. Online filtering stabilizes training by preserving informative samples while discarding low-quality ones.
5.3 Experience replay
Disabling experience replay causes a sharp performance drop, even below the vanilla GRPO baseline (Fig. 5). This counterintuitive result shows that high-quality samples from additional rollouts can be harmful without a replay mechanism.
@@ 5.3 Experience replay
Although high-quality samples drive high-entropy exploration, experience replay preserves their gradients, restoring stability and performance.
High-quality samples usually have low generation probabilities: rewarding them shifts responses toward rare but correct outputs, raising entropy, and disturbing convergence. Their large advantage yet low probability also makes their gradients prone to clipping, leaving only fleeting signals that are quickly overwritten. Without replay, this transient effect amplifies uncertainty and destabilizes training. Experience replay ensures these rare samples repeatedly contribute effective gradients, guiding convergence along a stable trajectory.
6 Conclusion
We examined the limitations of GRPO on complex reasoning tasks and identified two major issues: limited exploration and training instability. To address these, we proposed EFRame, an Exploration-Filter-Replay framework that enriches exploration with additional rollouts, stabilizes training via online filtering, and improves convergence by replaying rare but informative samples. This cycle transforms RL from a “reinforcement lottery” into a more structured and controllable process.
Ablation studies show that high-advantage positives are crucial for deeper reasoning, while low-advantage negatives drive premature convergence, highlighting the value of selective filtering. EFRame further enables fine-grained entropy control by jointly tuning rollout temperature and replay ratios, avoiding biases of entropy regularization and mitigating collapse. Future work will focus on adaptive balancing of exploration and replay, more principled filtering, and integration with verifiable rewards and workflow-based training, paving a path toward more reliable and scalable reasoning in LLMs.
Declaration of AI
AI is only used for translation and language polishing in this paper.
@@ 6 Conclusion
Appendix A Implementation details
We initialize all policy models and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$. GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward, while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens. For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
Appendix B Limitation
Although EFRame demonstrates substantial improvements across diverse benchmarks, it imposes a non-trivial requirement on the difficulty level of the training dataset. Specifically, EFRame benefits most when the dataset maintains a moderate difficulty level. If the dataset is overly challenging, both the regular rollout and the additional rollout stages suffer from excessively high failure rates, resulting in sparse reward signals. Conversely, if the dataset is too simple, the proportion of difficult prompts is low, rendering the additional rollout phase largely ineffective due to the scarcity of hard problems.
In this work, the selected datasets—DAPO-17K, the multimodal open-domain dataset, and Geometry3K—satisfy this moderate-difficulty criterion, enabling the additional rollout and filtering mechanisms to operate effectively. However, when training on the MATH-12K dataset, EFRame failed to deliver noticeable gains, with performance being nearly indistinguishable from the GRPO and DAPO baselines. This degradation is attributed to the dataset’s excessive simplicity, which results in an insufficient number of high-quality samples discovered during additional rollout exploration. This observation underscores that the success of EFRame is contingent on the availability of a sufficient proportion of moderately difficult samples to balance exploration efficacy and training stability.